OpenStudy (anonymous):

evaluate the limit:

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{ x ^{2} + 3x -2 }{3x ^{2} +4x -1 }\]

OpenStudy (anonymous):

Looks like a case where L'Hopital's rule will be useful. Always try to evaluate it straight up first. If we try that, we'll just get \[ \frac{\infty}{\infty} \] So let's try L'Hopital. Take the derivative of the numerator and denomerator, independently. This will give us \[\lim_{x->\infty} \frac{2x+3}{6x+4}\] Onece again, try to evaluate. We'll still get \[ \frac{\infty}{\infty} \] Use L'Hopital again: \[\lim_{x->\infty} \frac{2}{6}\] Alas! Evaluate the limit now to get \[\lim_{x->\infty} \frac{2}{6} = \frac{2}{6} = \frac{1}{3} \]

OpenStudy (anonymous):

A perhaps more clever way of doing this is to divide the entire limit piece by x^2, so we'd have \[ \lim_{ x-> \infty} \frac{1 + \frac{3}{x} - \frac{2}{x^2}}{3 + \frac{4}{x} -\frac{1}{x^2}} \] Since \[ \lim_{ x-> \infty} \frac{\text{integer}}{x} = 0\] and \[ \lim_{ x-> \infty} \frac{\text{integer}}{x^2} = 0\] We easily see that the limit is 1/3.

OpenStudy (anonymous):

yeah i see thanks man .. u help alot .. thanks